OpenStudy (anonymous):

Write the equation of the line through (2,0) with a slope of -3.

OpenStudy (ash2326):

Slope intercept form of a line \[y=mx+c\] m= slope of the line c=intercept @Katie.bily What's m here?

Parth (parthkohli):

We know that the equation of a line is y = mx + b. m is the slope. Now, we see that it is y = -3x + b Let's now substitute the given points. 0 = -3(2) + b 0 = -6 + b b = 6 Now, we know, that b = 6. Can you frame it now?